# UniHENN: Designing More Versatile Homomorphic Encryption-based CNNs without im2col

### v1

@@ Abstract
Homomorphic encryption enables computations on encrypted data without decryption, which is crucial for privacy-preserving cloud services. However, deploying convolutional neural networks (CNNs) with homomorphic encryption encounters significant challenges, particularly in converting input data into a two-dimensional matrix for convolution, typically achieved using the im2col technique. While efficient, this method limits the variety of deployable CNN models due to compatibility constraints with the encrypted data structure. UniHENN, a homomorphic encryption-based CNN architecture, eliminates the need for im2col, ensuring compatibility with a diverse range of CNN models using homomorphic encryption. Our experiments demonstrate that UniHENN surpasses the leading 2D CNN inference architecture, PyCrCNN, in inference time, as evidenced by its performance on the LeNet-1 dataset, where it averages 30.090 seconds–significantly faster than PyCrCNN’s 794.064 seconds. Furthermore, UniHENN outperforms TenSEAL, which employs im2col, in processing concurrent images, an essential feature for high-demand cloud applications. The versatility of UniHENN is proven across various CNN architectures, including 1D and six different 2D CNNs, highlighting its flexibility and efficiency. These qualities establish UniHENN as a promising solution for privacy-preserving, cloud-based CNN services, addressing the increasing demand for scalable, secure, and efficient deep learning in cloud computing environments.
@@
homomorphic encryption, privacy-preserving machine learning, convolutional neural networks, data privacy
@@ I Introduction
The growing use of machine learning (ML) in diverse domains such as finance, autonomous driving, industrial automation, computer vision, speech recognition, and healthcare has raised concerns regarding data privacy and security [35, 50]. These concerns are particularly relevant for cloud-based ML services, where sensitive data is processed on remote servers [3].
@@
Homomorphic encryption (HE) is a form of encryption that allows some computations on ciphertext without decrypting it. Therefore, it can be used as a powerful tool for preserving privacy in ML applications [22, 23, 20, 33, 13, 24, 30, 31, 11], enabling the confidentiality of sensitive information even when processed by a third party. HE is particularly relevant for cloud-based ML services, where sensitive data is often transmitted to remote servers for processing. With HE, cloud providers can ensure data privacy in their possession. For instance, a bank could employ HE to build an ML model for detecting fraudulent transactions without accessing customer data. Similarly, a healthcare provider could analyze patients’ medical records using HE without disclosing the patient’s identity. This potential has attracted interest from cloud service providers (e.g., Amazon AWS, Microsoft Azure, Naver Cloud) in integrating HE into ML-based cloud services, such as Data Box Frame [41].
@@
While convolution operations are fundamental to convolutional neural networks (CNNs), they demand substantial computational resources when performed on ciphertexts within an HE framework. Each kernel shift requires element-wise multiplications and additions, resource-intensive operations on ciphertexts. To reduce the computational load, most HE-based CNN implementations like TenSEAL [7] use the im2col function. This function transforms the input data into a matrix without discarding or altering the original information, thus allowing for more efficient computation. TenSEAL is applied in various research areas, including healthcare [21], federated learning [46], and biometrics [49]. However, this approach limits the versatility of CNN models to configurations that accommodate only a single convolution layer, potentially hindering their use in more complex structures. In practical scenarios, encrypted user data could be useful across multiple ML models. For instance, an encrypted patient image stored in a hospital database could be utilized for future analysis by various ML algorithms. Therefore, encrypting data without restricting it to a specific model would offer greater flexibility and utility.
In this paper, we introduce UniHENN, a privacy-preserving CNN model built using HE. UniHENN enables efficient CNN inference independent of the im2col function, resulting in encrypted input data that is usable across other ML models. We devise an effective computation method for convolution operations on flattened input within UniHENN, offering flexible reuse across various ML models. Furthermore, UniHENN supports batch operations, thereby facilitating the efficient processing of multiple ciphertexts. Our key contributions are summarized as follows:
@@
  • We introduce UniHENN, a novel and efficient CNN model inference mechanism based on HE, which facilitates input ciphertext reusability. Unlike other CNN implementations that require a specific model input structure for the im2col function, UniHENN is designed to handle model-free input ciphertexts, enabling the use of encrypted input data across various HE-based ML services without the need for re-encryption. We demonstrate the effectiveness of this approach by successfully constructing and training seven different CNN models on four datasets: MNIST [15], CIFAR-10 [26], USPS [19], and electrocardiogram (ECG) [38]. The source code for UniHENN is available at https://github.com/hm-choi/uni-henn.
@@
  • We empirically demonstrate that UniHENN is efficient. Experimental results indicate an average inference time of 30.090 seconds, significantly outperforming the state-of-the-art HE-based 2D CNN inference architecture PyCrCNN [16], which requires an average of 794.064 seconds for inference.
@@
  • We introduce a batch processing strategy for UniHENN to handle multiple data instances in a single operation. This strategy efficiently combines multiple data instances into a single ciphertext, reducing the inference time for 10 MNIST images to 16.248 seconds. It outperforms TenSEAL’s CNN model [7], which takes 63.706 seconds. While UniHENN is less efficient than TenSEAL for processing a single image, it surpasses TenSEAL’s performance when processing $k$ images simultaneously, where $k\geq 3$.
@@ II-A Convolutional Neural Network (CNN)
Convolutional neural networks (CNNs) were first introduced by LeCun et al. [28] for processing grid-structured data, such as images. The LeNet-1 model, presented in 1989 [27], comprised two convolutional layers, two average pooling layers, and one fully connected layer, marking the first concept of the LeNet architecture.
In 1998, the LeNet-5 model was introduced [29], consisting of three convolutional layers, two pooling layers, and two fully connected layers. It was designed to recognize handwritten postal codes efficiently, outperforming traditional methods like the multilayer perceptron model.
CNNs gained significant traction in the 2010s, primarily due to their exceptional performance on large datasets like ImageNet [14], which includes millions of images across 1,000 classes. Currently, research into the utilization of CNNs, such as DGCNN and its applications, is ongoing [47].
A typical CNN includes an input layer, multiple hidden layers, and an output layer. The input layer receives the image and forwards its pixel values into the network. Since images are generally matrices, the node count in the input layer corresponds to the image size. The hidden layers, situated between the input and output layers, extract relevant features using convolution, activation, and pooling operations. Each hidden layer processes information from the input or preceding layers, enabling iterative learning. This information is then relayed to subsequent layers, culminating in the final prediction or classification at the output layer.
Kiranyaz et al. [25] introduced a 1D CNN for disease-specific ECG classification. A 1D CNN, tailored for one-dimensional data, is particularly suitable for signal data, time-series data, and text data. In this architecture, both the input and the convolution filter are one-dimensional; the filter slides over the input in the convolution layer to perform operations.
@@ II-B Homomorphic Encryption (HE)
Before encrypting with CKKS, it is crucial to create a plaintext structure that mimics the input data. The number of slots specifies the vector size that can be encrypted and is set during parameter selection. CKKS supports three basic operations: addition, multiplication of ciphertexts, and rotation of ciphertext.
Let $N$ be the total degree of the CKKS parameter then the number of slots is $N/2$. Two real vectors $\mathbf{v_{1}},\mathbf{v_{2}}$ and $C(\mathbf{v_{1}}),C(\mathbf{v_{2}})$ denotes the ciphertext of the vector $\mathbf{v_{1}},\mathbf{v_{2}}$. (P) means that the operation is defined between a ciphertext and a plaintext, and (C) means that the operation is defined between two ciphertexts. Then the operations, addition ($Add$), multiplication ($Mul$), and rotation $(Rot)$ can be represented as follows:
@@
  • Rotation :
    $Rot(C(\mathbf{v}),r)=C(v_{r},v_{r+1},...,v_{N/2-1},v_{0},...,v_{r-1})$
    where $\mathbf{v}=(v_{1},v_{2},...,v_{N/2})$ and $r$ is a positive integer.
The notation +,×+$\times$+ , × in the plaintext addition and multiplication are elementwisely defined. The depth specifies the maximum number of multiplications, which affects the noise level in the ciphertext, and exceeding this limit may compromise decryption. The level specifies the allowed number of the multiplication. The depth is fixed when the parameter is determined but the level can be decreased when multiplications are conducted.
For practical HE implementations, various libraries are available, including SEAL-Python [2], Lattigo [39], HElib [18], and OpenFHE [5]. We selected SEAL-Python for its efficient support of the CKKS scheme.
II-C HE Libraries
Several libraries have been developed in the HE field to facilitate practical implementations. Among these, SEAL-Python [44], Lattigo [39], HElib [1], and OpenFHE [5] are popularly used.
SEAL [44] supports various HE schemes (BFV and CKKS) and incorporates optimizations to speed up homomorphic operations, such as ciphertext packing and batching. For UniHENN, we chose SEAL-Python, a Python adaptation of the original C++ SEAL library. We selected SEAL-Python for its user-friendly interface. However, UniHENN could also be implemented with other libraries.
Lattigo [39], written in Go, focuses on ring-learning-with-errors-based HE primitives and multiparty-HE-based secure protocols. Its cross-platform compatibility and optimized protocols enable fast and secure computation.
HElib [1] specializes in the BGV and CKKS schemes, incorporating bootstrapping and several optimizations to speed up homomorphic evaluation. It effectively uses the Smart-Vercauteren ciphertext packing techniques and the Gentry-Halevi-Smart optimization. Also, it provides an assembly language for HE, which includes low-level routines, multithreading capabilities, and more.
OpenFHE [5], supporting both bootstrapping and various hardware acceleration backends, utilizes a standard hardware abstraction layer to ensure compatibility and efficiency.
@@ III Overview of UniHENN
    UniHENN uses average pooling, eliminating multiplication. This configuration allows for more layers within given parameter settings, giving service providers more flexibility to incorporate average pooling without concern for operation count.
@@
Figure 1 presents a high-level overview of UniHENN’s operational flow. The process begins when a client encrypts one or more images using a public key. These encrypted images are consolidated into a single ciphertext and sent to a cloud service specializing in data analytics. Each service (i.e., Service 1, Service 2, …, Service N) uses its unique CNN model for data processing, executing specific algorithms and computations on the received ciphertext with the evaluation key. Each CNN model has a distinct layer architecture and optimized parameters. After processing the ciphertext through their respective CNN models, the encrypted inference results are returned to the client. The client decrypts these results using the corresponding secret key to obtain the processed outcomes from each CNN model. This framework allows the client to benefit from multiple CNN models while preserving the confidentiality of the data, indicating a significant step forward in privacy-preserving machine learning.
In Table I, we summarize the important notations used in this paper with UniHENN.
@@
Notation
Definition
$N$
Number of slots of ciphertext in power of 2.
$D$
Depth (or level) of ciphertext, representing the current limit on the number of multiplication operations allowed for the ciphertext.
$CH_{in},CH_{out}$
Number of input and output channels in a layer.
$DAT_{in},DAT_{out}$
Number of input and output data in a fully connected (FC) layer.
$R_{q}$
Polynomial quotient ring $(\mathbb{Z}_{q}/\langle X^{N}+1\rangle)$, used to create a vector space for the ciphertext.
$W_{img},H_{img}$
Width and height of the used image data.
$K$
Width and height of the kernel.
$W_{in},H_{in}$
Width and height of input data for this layer.
$W_{out},H_{out}$
Width and height of output data for this layer.
$S_{total}$
Used in Section IV-E. The value is the product of each convolutional layer’s stride and the kernel sizes of all average pooling layers.
@@ III-A Construction of Input Data
Figure 2 illustrates an example of data transformation for encryption. The input data consists of $W_{img}\times H_{img}$ numbers. We flatten the original data row by row, as shown in Figure 2. Implementing this flattening procedure ensures universal compatibility of the input data with ML models, enhancing its adaptability to the algorithms used by ML services.
@@ III-B Combining Ciphertexts for Batch Operation
For example, as illustrated in Figure 3, if 800 slots are needed to process the input data through the CNN model, then the $i$-th input data would be rotated to the right by $(i-1)\times 800$ slots, and the rotated vector is added to the ciphertext. UniHENN can integrate any number of ciphertexts into a single one, provided the total size of the encrypted data does not exceed the number of slots. This method enhances the efficiency of UniHENN when handling multiple ciphertexts concurrently.
@@ IV-A Construction of the Drop-DEPTH
The parameter is pre-defined because UniHENN’s design philosophy is to support the ciphertext that is encrypted before the model is fixed. So, the CKKS parameter is determined and encrypted before the CNN model is defined for reusability of the input ciphertext independent of the model. But, the multiplication and rotation time of HE is linearly increased dependent on the depth. Thus, the best choice of the CKKS parameter is determined by the number and structure of the CNN model’s layers. However, in this scenario, the depth is defined before when the CNN model is fixed, which means that the depth can be larger than the exact number of total levels of the CNN’s model. Thus, we suggest the Drop-DEPTH, a novel approach to reduce the depth that is optimized when the target CNN model. If the input ciphertext has depth $D$ and the target CNN model’s level is $L$ where $D\leq L$ then the Drop-DEPTH reduces the depth of the input ciphertext as $L$ with multiply $D-L$ ciphertexts that are the encryption of vectors where all elements are consists of $1$. By these simple structure, time of execution is very fast (In our experiments in Section VII, the time of Drop-DEPTH operation does not exceed $100ms$.

### v3

@@ Abstract
Homomorphic encryption (HE) enables privacy-preserving deep learning by allowing computations on encrypted data without decryption. However, deploying convolutional neural networks (CNNs) with HE is challenging due to the need to convert input data into a two-dimensional matrix for convolution using the im2col technique, which rearranges the input for efficient computation. This restricts the types of CNN models that can be used since the encrypted data structure must be compatible with the specific model. UniHENN is a novel HE-based CNN architecture that eliminates the need for im2col, enhancing its versatility and compatibility with a broader range of CNN models. UniHENN flattens input data to one dimension without using im2col. The kernel performs convolutions by traversing the image, using incremental rotations and structured multiplication on the flattened input, with results spaced by the stride interval. Experimental results show that UniHENN significantly outperforms the state-of-the-art 2D CNN inference architecture named PyCrCNN in terms of inference time. For example, on the LeNet-1 model, UniHENN achieves an average inference time of 30.089 seconds, about 26.6 times faster than PyCrCNN’s 800.591 seconds. Furthermore, UniHENN outperforms TenSEAL, an im2col-optimized CNN model, in concurrent image processing. For ten samples, UniHENN (16.247 seconds) was about 3.9 times faster than TenSEAL (63.706 seconds), owing to its support for batch processing of up to 10 samples. We demonstrate UniHENN’s adaptability to various CNN architectures, including a 1D CNN and six 2D CNNs, highlighting its flexibility and efficiency for privacy-preserving cloud-based CNN services.
@@
homomorphic encryption, privacy-preserving machine learning, data privacy
[…]
=-15pt
@@ I Introduction
The widespread adoption of deep learning has expanded beyond traditional tasks like image classification [50] to diverse applications such as attack detection [47] and resource extraction analysis [49], with machine learning (ML) now playing a crucial role in sensitive sectors including finance, healthcare, and autonomous systems. However, this proliferation has raised significant privacy concerns [36, 53], particularly for cloud-based ML services processing sensitive data on remote servers [4]. Balancing ML’s powerful capabilities with the need to safeguard individual privacy has become a critical challenge, highlighting the urgent need for advanced privacy-preserving techniques in ML applications handling sensitive information.
@@
Homomorphic encryption (HE) is a powerful tool for preserving privacy in ML applications [25, 22, 34, 14, 27, 26]. For example, Choi et al. introduced privacy-preserving biometric authentication systems [12, 11] that leverage HE for web and cloud environments. HE allows computations to be performed on encrypted data without decryption, ensuring that sensitive information remains confidential even when processed by a third party. This is particularly beneficial for cloud-based ML services, where sensitive data is often transmitted to remote servers for processing. By using HE, cloud providers can ensure data privacy, such as a bank training an ML model to detect fraudulent transactions without accessing customer data or a healthcare provider analyzing patient medical records without disclosing patient identities. Additionally, faster inference time with privacy preservation is crucial for real-time applications in these sensitive fields, where quick and secure data processing can significantly improve decision-making and user experience.
@@
Convolution operations are essential to Convolutional Neural Networks (CNNs) but require significant computational resources when performed on ciphertexts in HE. Each kernel shift requires element-wise multiplications and additions, resource-intensive operations on ciphertexts. To reduce the computational load, most HE-based CNN implementations like TenSEAL [7] use the im2col function. This function transforms the input data into a matrix without discarding or altering the original information, thus allowing for more efficient computation. TenSEAL is applied in various research areas, including healthcare [24], federated learning [48], and biometrics [51]. However, this approach limits the versatility of CNN models to configurations that accommodate only a single convolution layer, potentially hindering their use in more complex structures. In practical scenarios, encrypted user data could be useful across multiple ML models. For instance, an encrypted patient image stored in a hospital database could be utilized for future analysis by various ML algorithms. Therefore, encrypting data without restricting it to a specific model would offer greater flexibility and utility.
We introduce UniHENN, a privacy-preserving CNN model using HE that enables efficient CNN inference without relying on the im2col function. Our novel convolution algorithm flattens input data into a one-dimensional form, eliminating the need for data rearrangement. The kernel traverses the input, performing convolutions with incremental rotations and structured multiplication, ensuring only relevant elements are used. This approach overcomes im2col limitations by reducing image size dependency and focusing on kernel size, improving efficiency and flexibility across various ML models. UniHENN also supports batch operations for efficient multi-ciphertext processing. Our key contributions include:
@@
  • We introduce UniHENN, a novel CNN model inference mechanism based on HE, which facilitates input ciphertext reusability. Unlike other CNN implementations that require a specific model input structure for the im2col function, UniHENN is designed to handle model-free input ciphertexts, enabling the use of encrypted input data across various HE-based ML services without the need for re-encryption. We demonstrate the effectiveness of this approach by successfully constructing and training seven different CNN models on four datasets: MNIST [16], CIFAR-10 [29], USPS [21], and electrocardiogram (ECG) [39]. The source code for UniHENN is available at https://github.com/hm-choi/uni-henn.
@@
  • We empirically demonstrate the efficiency of UniHENN. Experimental results indicate an average inference time of 30.089 seconds, significantly outperforming the state-of-the-art HE-based 2D CNN inference architecture PyCrCNN [17], which requires an average of 800.591 seconds for inference.
@@
  • We introduce a batch processing strategy for UniHENN to handle multiple data instances in a single operation. This strategy efficiently combines multiple data instances into a single ciphertext, reducing the inference time for 10 MNIST images to 16.247 seconds. It outperforms TenSEAL’s CNN model [7], which takes 63.706 seconds. While UniHENN is less efficient than TenSEAL for processing a single image, it surpasses TenSEAL’s performance when processing multiple images simultaneously, particularly when $k\geq 3$.
@@ II-A Convolutional Neural Network (CNN)
Convolutional Neural Networks (CNNs) were first introduced by LeCun et al. [31] for processing grid-structured data like images. In 1989, the LeNet-1 model was presented [30], comprising two convolutional layers, two average pooling layers, and one fully connected layer. Which is the first concept of the LeNet architecture.
In 1998, the LeNet-5 model was presented [32], comprising three convolutional layers, two pooling layers, and two fully connected layers. It was designed to efficiently recognize handwritten postal codes, outperforming traditional methods like the multilayer-perceptron model.
CNNs gained significant traction in the 2010s, primarily due to their excellent performance on large datasets like ImageNet [15], which includes millions of images across 1,000 classes.
A typical CNN consists of an input layer, multiple hidden layers, and an output layer. The input layer receives the image and forwards its pixel values into the network. Since images are generally matrices, the node count in the input layer corresponds to the image size. The hidden layers, situated between the input and output layers, extract relevant features using convolution, activation, and pooling operations. Each hidden layer receives information from either the input or preceding layers, facilitating iterative learning. This information is then passed to subsequent layers, culminating in the final prediction or classification at the output layer.
Kiranyaz et al. [28] introduced a 1D CNN for disease-specific ECG classification. A 1D CNN is a variant of CNNs tailored for one-dimensional data, making it particularly suitable for signal data, time-series data, and text data. In this architecture, both the input and the convolution filter are one-dimensional; the filter slides over the input in the convolution layer to execute operations.
@@ II-B Homomorphic Encryption (HE)
CKKS encryption requires structuring plaintext to mirror input data. The number of slots, defined during parameter selection, determines the encryptable vector size. CKKS supports three fundamental ciphertext operations: addition, multiplication, and rotation.
Let $N$ be the total degree of the CKKS parameter, then the number of slots is $N/2$. Two real vectors $\mathbf{v_{1}},\mathbf{v_{2}}$ and $C(\mathbf{v_{1}}),C(\mathbf{v_{2}})$ denote the ciphertext of the vectors $\mathbf{v_{1}},\mathbf{v_{2}}$. (P) indicates that the operation is defined between a ciphertext and a plaintext, and (C) indicates that the operation is defined between two ciphertexts. Then the operations, addition ($Add$), multiplication ($Mul$), and rotation $(Rot)$ can be represented as follows:
@@
  • Rotation: $,Rot(C(\mathbf{v}),r)=C(v_{r},v_{r+1},\ldots,v_{N/2-1},v_{0},$$\ldots,v_{r-1})$, where $\mathbf{v}=(v_{0},v_{1},\ldots,v_{N/2-1})$ and $r$ is a positive integer.
The operations $+$ and $\times$ represent elementwise addition and multiplication in the plaintext space, respectively. The depth parameter specifies the maximum number of sequential multiplications that can be performed while ensuring correct decryption, as each multiplication operation increases ciphertext noise. Exceeding this limit may result in decryption failure. While depth is predetermined during parameter selection, level dynamically indicates the remaining multiplication capacity, decreasing with each multiplication operation executed.
For practical HE implementations, several general-purpose HE libraries are available, including SEAL-Python [2], Lattigo [40], HElib [20], and OpenFHE [5]. We selected SEAL-Python for its efficient support of the CKKS scheme. Building on this foundation, we developed UniHENN, a HE-based framework specifically optimized for CNN inference. While currently implemented with SEAL-Python, UniHENN is designed to be adaptable to other libraries.
@@ III Overview of UniHENN
    UniHENN opts for average pooling, which eliminates the need for multiplication, thereby reducing operation time. This configuration allows for more layers within given parameter settings, giving service providers more flexibility to incorporate average pooling without concern for operation count.
@@
Figure 1 presents a high-level overview of UniHENN’s operational flow. The process begins when a client encrypts one or more images using a public key. These encrypted images are consolidated into a single ciphertext and sent to a cloud service specializing in data analytics. Each service (i.e., Service 1, Service 2, …, Service $K$) uses its unique CNN model for data processing, executing specific algorithms and computations on the received ciphertext with the evaluation key. Each CNN model has a distinct layer architecture and optimized parameters. After processing the ciphertext through their respective CNN models, the encrypted inference results are returned to the client. The client decrypts these results using the corresponding secret key to obtain the processed outcomes from each CNN model. This framework allows the client to benefit from multiple CNN models while preserving the confidentiality of the data, indicating a significant step forward in privacy-preserving machine learning. Table I summarizes the notations used throughout this paper.
@@
Notation Definition
$N$ Degree of polynomial. It is in the form of a power of two.
$D,L$ Depth and level of ciphertext. This represents the current remaining limit on the number of multiplication operations allowed for the ciphertext.
$CH_{in},CH_{out}$ Number of input and output channels in layer
$DAT_{in},DAT_{out}$ Number of input and output data in FC layer
$R_{q}$ Polynomial quotient ring $(\mathbb{Z}_{q}/\langle X^{N}+1\rangle)$. It is used to create a vector space of the ciphertext.
$W_{img},H_{img}$ Width and height of the input image data
$K$ Width and height of the kernel
$W_{in},H_{in}$ Width and height of input data from this layer
$W_{out},H_{out}$ Width and height of output data from this layer
$S_{total}$ It is used in Section IV-E. The value is the multiplication of each convolutional layer’s stride and the kernel sizes of all average pooling layers.
@@ III-A Construction of Input Data
Figure 2 shows an example of data transformation for encryption. The input data consists of $W_{img}\times H_{img}$ numbers. We flatten the original data row by row, as shown in Figure 2. This flattening procedure ensures universal compatibility of the input data with ML models, enhancing adaptability to algorithms used by ML services.
@@ III-B Combining Ciphertexts for Batch Operation
Figure 3 illustrates one example. In this case, 800 slots are needed to process the input data through the CNN model. Thus, the $i$-th input data would be rotated to the right by $(i-1)\times 800$ slots, and the rotated vector is added to the ciphertext. UniHENN can integrate any number of ciphertexts into a single one, provided the total size of the encrypted data does not exceed the number of slots. This method enhances the efficiency of UniHENN when handling multiple ciphertexts concurrently.
@@ IV-A Construction of the Drop-DEPTH
The parameter is pre-defined because UniHENN’s design philosophy is to support the ciphertext that is encrypted before the model is fixed. So, the CKKS parameter is determined and encrypted before the CNN model is defined for reusability of the input ciphertext independent of the model. But, the multiplication and rotation time of HE is increased dependent on the level. Thus, the best choice of the CKKS parameter is determined by the number and structure of the CNN model’s layers. However, in this scenario, the depth is defined before when the CNN model is fixed, which means that the depth can be larger than the exact number of total levels of the CNN’s model. Thus, we suggest the Drop-Level, a novel approach to reduce the level that is optimized when the target CNN model. If the input ciphertext has depth $D$ and the target CNN model’s level is $L$ where $L\leq D$ then the Drop-Level reduces the level of the input ciphertext as $L$ with multiply $D-L$ ciphertexts that are the encryption of vectors where all elements are consists of $1$. By these simple structures, the time of execution is very fast (In our experiments in Section VII, the time of Drop-Level operation does not exceed $100ms$.